Your company can produce and sell only one of the following two products;

Machine Hours Required

Contribution Margin

Product 1

6

SR 60

Product 2

4

SR 50

If the company has machine capacity of 1,500 hours, the total contribution margin of the product selected should maximize net income produced to SR 16,000.

Opinion

Justification
Answer #1

Contribution margin per hour

Product 1 = 60/6 = 10 per hour

Product 2 = 50/4 = 12.5 per hour

Opinion:

As per the results product 2 should be produced as its contribution margin per hour is higher than product 1.

Justification:

The total contribution margin of product 2 is 18,750 (1500*12.5) which exceeds limit of 16,000 as compared to product 1 total contribution margin 15,000 (1500*10) which is less than the limit of 16,000.